OpenStudy (eric_d):

If y=1/2( e^x + e^-x ), show that (dy/dx)^2=y^2-1

OpenStudy (eric_d):

@Jhannybean

Miracrown (miracrown):

(dy/dx)^2 doesn't mean 2nd derivative there, right?

OpenStudy (jhannybean):

implicitly differentiate both sides.

Miracrown (miracrown):

So dy/dx by itself means the 1st derivative So (dy/dx)^2 means the first derivative squared

Miracrown (miracrown):

So you need to find the derivative of y = (1/2)(e^x + e^(-x)) and square it Then simplify to show that (dy/dx)^2 is the same as y^2 - 1

ganeshie8 (ganeshie8):

if you know hyperbolics @eric_d ...

OpenStudy (eric_d):

Nope

Miracrown (miracrown):

If you're looking for an easier method than what I have stated above, then you have to compute the derivative either way since you have dy/dx on the left

OpenStudy (jhannybean):

\[\sinh = \frac{e^x-e^{-x}}{2}\]

OpenStudy (eric_d):

why e^x-e^-x / 2 instead of (e^x + e^-x) / 2 ?

Miracrown (miracrown):

You can use hyperbolic trig functions though y = (1/2)(e^x + e^(-x)) = cosh(x) But you would still need to take the derivative

OpenStudy (jhannybean):

Oh, whoops.cosine is the positive one, thought it was sine.

OpenStudy (jhannybean):

you can just as easily take the derivative of each part.

Miracrown (miracrown):

That's what I said ...

ganeshie8 (ganeshie8):

he was not introduced to hyperbolics yet.. so may be lets continue with e^x...

OpenStudy (jhannybean):

\[\frac{d}{dx}(y) = \frac{1}{2} \left[(e^x)' +(e^{-x})'\right]\]

Miracrown (miracrown):